OpenStudy (anonymous):

Is the following system consistent or inconsistent? 8x + y + 4 = 0 y = -8x - 4

OpenStudy (superdavesuper):

Leave y on the left hand of the 1st eqn n move everything else to the right, how the 1st eqn looks?

OpenStudy (anonymous):

8x+y=-4 +8x-y =-4 16x=-8 16/-8=-2 x=-2

OpenStudy (superdavesuper):

Think u make a typo somewhere - plz check again? :)

OpenStudy (anonymous):

Its consistent because I can solve it ?

OpenStudy (superdavesuper):

Yes on consistent but not because u can solve it....